# Mixed-Integer Optimal Control via Reinforcement Learning: A Case Study on Hybrid Vehicle Energy Management

### v2

@@ 7       Compute discrete action $a_{t}^{d}$ according to $\varepsilon$-greedy
$\begin{cases}&\mathrm{Pr}(a_{t}^{d}=\text{a sample randomly from }\mathcal{A}_{d})=\varepsilon,\\ &\mathrm{Pr}(a_{t}^{d}=\arg\max_{a^{d}}Q_{\omega_{1}}\left(s_{t},a^{d},a_{t}^{c}\right))=1-\varepsilon\end{cases}$
Take action $a_{t}=(a_{t}^{d},a_{t}^{c})$, observe reward $r_{t}$ and the next state $s_{t+1}$;
8       Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ in replay buffer $\mathcal{B}$;
9       Sample $B$ transitions $\{(s_{i},a_{i},r_{i},s_{i+1})\}_{i\in[B]}$ from $\mathcal{B}$;
10       Compute target continuous action
@@
Compute the target value
@@
Update the two Q-networks ($j=1,2$) by one step of gradient descent
$\omega_{j}\leftarrow\omega_{j}-\alpha\nabla_{\omega_{j}}\frac{1}{|B|}\sum_{i\in[B]}\left(Q_{\omega_{j}}\left(s_{i},a_{i}^{d},a_{i}^{c}\right)-y_{i}\right)^{2}$
if t mod $K$ == 0 then
11             Update the actor network by one step of gradient ascent
$\theta\leftarrow\theta+\beta\nabla_{\theta}\frac{1}{|B|}\sum_{i\in[B]}\max_{a^{d}}Q_{\omega_{1}}\left(s_{i},a^{d},\mu_{\theta}\right)$
Update the target networks ($j=1,2$) with
$\omega_{j}^{\prime}\leftarrow\rho\omega_{j}+(1-\rho)\omega_{j}^{\prime},\quad\theta^{\prime}\leftarrow\rho\theta+(1-\rho)\theta^{\prime}$
@@ 11             Update the actor network by one step of gradient ascent
III PHEV Energy Management Problem Formulation
@@
III-A PHEV Powertrain Configuration
@@
This study focuses on a parallel hybrid electric vehicle [27]. The powertrain architecture comprises an internal combustion engine, an electric motor, a clutch, and an automated manual transmission, as depicted in Fig. 2. The vehicle’s HEV mode can be altered by engaging or disengaging the clutch between the engine and the motor, providing three modes of operation: motor-only, engine-only, and hybrid-drive.
III-B Vehicle Dynamics
The vehicle dynamics are described by the following equations:
$T_{w}=\left[ma+\frac{1}{2}C_{d}\rho Av^{2}+\mu mg\cos\theta+mg\sin\theta\right]r$ (14)
@@ III-B Vehicle Dynamics
where $T_{w}$ is the wheel torque, $m$ is the vehicle mass, $a$ is the vehicle acceleration, $C_{d}$ is the air drag coefficient, $\rho$ is the air density, $A$ is the vehicle cross section area, $\omega$ is the shaft angular velocity, $v$ is the vehicle velocity, $\mu$ is the rolling resistance coefficient, $g$ is the gravitational acceleration, $\theta$ is the road grade, $r$ is the tyre radius, $i_{d}$ is the final drive gear ratio, and $i_{g}$ is the transmission gear ratio.
@@
where $T_{e,d}$ and $T_{m}$ represent the driving torque transmitted from the engine and motor to the shaft, respectively. $c$ is the clutch state with $c=1$ representing the clutch engaged and $c=0$ representing the clutch disengaged. $\eta_{d}$ is the final drive efficiency, $\eta_{g}$ is the transmission efficiency, and $T_{b}$ is the mechanical brake torque.
III-C Engine Model
@@
The engine torque is separated into two parts: the driving torque $T_{e,d}$ and the idle torque $T_{e,i}$, which can be expressed as:
$T_{e}=T_{e,d}+T_{e,i}$ (17)
@@
where $T_{e,max}=f_{e}(\omega_{e})$ represents the maximum engine torque at a given angular velocity $\omega_{e}$, which is depicted in Fig. 3. When the clutch is disengaged, $T_{e,d}$ is set to zero, and $T_{e}=T_{e,i}$. The idle torque primarily serves to power the auxiliary systems. The engine speed is governed by the following equation:
@@
where $\omega_{m}$ is the motor angular velocity, and $\omega_{ei}$ is the engine idle angular velocity. The constraint on $\omega_{e}$ is imposed to ensure that the engine speed does not drop below the idle speed when the clutch is engaged.
@@ III-D Motor Model
III-E Battery Model
Figure 5: The open-circuit voltage and resistance of the battery cell.
We consider a lithium-ion battery pack consisting of 112 cells, where the cells are connected in series. The battery model is based on the battery state of charge (SOC) and the battery open-circuit voltage $E$. The battery SOC is calculated by:
$\dot{SOC}=-\frac{E-\sqrt{E^{2}-4R_{b}P_{b}}}{2R_{b}Q_{\max}}$ (25)
$P_{b}=T_{m}w\eta_{m}^{\operatorname{sgn}\left(-T_{m}\right)}$ (26)
@@ III-E Battery Model
where $I_{b}$ is the net current, $Q_{max}$ is the battery capacity, $R_{b}$ is the battery resistance, $P_{b}$ is the battery power, $\eta_{m}$ is the motor efficiency, and $f_{\eta}$ is the motor efficiency interpolation function, which is shown in Fig. 4. The battery open-circuit voltage and resistance are denoted as functions of the battery SOC, as illustrated in Fig. 5.
III-F Problem Formulation
@@
where $k_{f}$ and $k_{b}$ are the fuel and electricity prices, respectively, $\tilde{p}$ is the reference penalty, which is set to be the intermediate value of fuel consumption, $p_{\varsigma}$ and $p_{c}$ are the penalties for gear shift and clutch engagement/disengagement, respectively, $\zeta$ and $\xi$ are the penalty coefficients, $\Delta t$ is the control interval, $T_{0}$ and $T_{cyc}$ are the initial and final time steps of the drive cycle, respectively. The drive cycle utilized for training is the China heavy-duty transient cycle for light trucks (CHTC-LT).
IV Experiments and Discussion
IV-A TD3AQ for PHEV Energy Management
IV-A1 State
@@ IV-A2 Action
where $T_{e}$ is continuous, and $\varsigma$ and $c$ are discrete, making it an MIOC problem. In negative demand torque, regenerative braking is prioritized for battery charging. If regenerative braking is insufficient, mechanical braking provides the remaining torque. In positive demand torque, the mechanical braking torque is set to zero, and the engine and motor supply the demand torque. Thus, the proposed action determines the vehicle’s torque distribution entirely.
@@
IV-A3 Reward
In addition to the cost function defined in Eq. (29), it is imperative to ensure that the system dynamics remain within specific constraints. Thus, the reward function is designed as a combination of fuel consumption and the penalty for violating the constraints. The detailed hand-crafted reward function is defined as follows:
$r=r_{cost}-p_{\omega}-p_{T_{m}}-p_{SOC}$ (34)
@@ IV-A3 Reward
IV-B Comparison Benchmarks
Four benchmark methods, DP, DDAQ [25], PATD3 [14], and Rainbow [11], are selected for comparison. DP employs the MATLAB toolbox developed by Sundström et al. [28]. The continuous engine torque is discretized with an interval of 25 Nm. Empirical experiments have demonstrated that finer discretization does not improve the performance significantly (less than 0.05% improvement in cost with 10 Nm discretization while requiring seven times more computation time). Since Rainbow is designed for discrete action spaces, it adopts the same discretization scheme as DP. All RL algorithms utilize Optuna [29] to tune the hyperparameters.
IV-C Experimental Setup
The training process involves a total of 200 000 timesteps, where each timestep represents one second of driving time. Model evaluation was performed every 10 000 timesteps, with each evaluation period lasting one episode, i.e., one drive cycle. The final model for testing was selected based on the highest return achieved during evaluation. All experiments were conducted on a desktop computer equipped with an AMD Ryzen 9 5950X CPU and an NVIDIA GeForce RTX 3080 Ti GPU. The source code of the algorithm is available at https://github.com/xujinming01/TD3AQ.
IV-D Optimal Control Performance
@@
(d) Clutch state (0 = disengaged, 1 = engaged)
Figure 6: The hybrid action output of DP and TD3AQ for CHTC-LT.
TABLE I: Simulation Results for CHTC-LT
@@
TD3AQ 116 56 0 0 0 15.66 4.13%
DDAQ 87 48 0 0 0 16.21 7.80%
PATD3 226 174 1 5 0 16.58 10.21%
Rainbow 251 208 0 1 0 16.22 7.85%
Fig. 6 displays the hybrid action output of DP and TD3AQ employed in CHTC-LT. It is evident that TD3AQ exhibits similar output trends to DP, however, with a higher frequency of gear shifts and clutch state changes. Notably, DP operates under the assumption that the entire drive cycle is known in advance. In contrast, RL algorithms only possess access to the current state, yielding suboptimal solutions. Table I presents the simulation results for CHTC-LT, where $\tau_{sh}$ and $\tau_{cl}$ denote the number of gear shifts and clutch engagements, respectively, $\nu_{T_{m}}$, $\nu_{\omega}$, and $\nu_{soc}$ denote the instances where motor torque, shaft angular velocity, and SOC exceed the predefined constraints. Among the RL algorithms, TD3AQ exhibits the lowest cost, with a gap of 4.13% compared to DP. Furthermore, TD3AQ and DDAQ do not violate any constraints, while PATD3 and Rainbow demonstrate a few violations in terms of motor torque and shaft angular velocity constraints.
@@
TABLE II: Wall-Clock Time Comparison for CHTC-LT
Strategy Training time (h) step time (ms)
TD3AQ 0.18 0.38
DDAQ 0.17 0.36
PATD3 0.18 0.74
Rainbow 0.59 0.97
DP N/A 96599
Table II presents the time consumption comparison for CHTC-LT, where step time represents the time required to execute a single action during testing. While DP can obtain the optimal solution for each step in a single calculation, it comes at the expense of a significant amount of time, with its single solution time being six orders of magnitude greater than RL. This substantial discrepancy in time consumption highlights the efficiency of the RL algorithms, as indicated by their step times falling within the millisecond range. Consequently, RL demonstrates promising potential for real-time optimal control applications.
IV-E Generalization Results
TABLE III: Generalization Results of the New Cycles
@@ IV-E Generalization Results
TD3AQ 1 0 0 14.04 3.23%
DDAQ 0 0 0 14.61 7.41%
PATD3 6 2 0 14.77 8.60%
Rainbow 0 0 0 14.76 8.48%
@@
TD3AQ 4 0 0 9.88 5.94%
DDAQ 0 8 0 10.55 13.18%
PATD3 14 7 0 9.97 6.95%
Rainbow 10 7 0 10.38 11.36%
@@
TD3AQ 0 0 0 5.72 3.63%
DDAQ 0 0 0 5.89 6.78%
PATD3 0 1 0 5.80 5.16%
Rainbow 1 0 0 6.13 10.99%
To evaluate the generalization capabilities of the proposed CDRL algorithm TD3AQ, the model trained exclusively on the CHTC-LT drive cycle is tested on three additional drive cycles: JE05, HD-UDDS, and WHVC. The simulation results are summarized in Table III.
The findings demonstrate that the TD3AQ algorithm exhibits commendable generalization performance when applied to new drive cycles. In comparison with other RL algorithms, TD3AQ closely approximates the cost achieved by the DP strategy, with an average gap of 4.27% across the three cycles. However, the remaining RL algorithms display a larger deviation in terms of cost.
@@
V Conclusion
In this brief, we present a novel reinforcement learning algorithm called TD3AQ for solving mixed-integer optimal control problems. The TD3AQ algorithm uses actor-Q networks to directly generate continuous and discrete actions from the current state. We evaluate the proposed TD3AQ algorithm on a hybrid electric vehicle energy management problem and show that it achieves satisfactory performance with a cost gap of 4.13% compared to the DP strategy. TD3AQ’s ability to generalize is demonstrated through the results on three new drive cycles. The algorithm’s sub-millisecond response time indicates its significant potential for real-time, mixed-integer optimal control applications. However, note that the algorithm carries the risk of violating constraints, which, if unaddressed, could potentially lead to vehicle and battery damage in real-world scenarios. Future work will focus on extending TD3AQ to more complex, higher-dimensional problems to further substantiate its adaptability and robustness, and developing CDRL algorithms with safety guarantees.

### v3

@@ 13      Compute discrete action $a_{t}^{d}$ according to $\varepsilon$-greedy
$a_{t}^{d}=\begin{cases}\text{a sample randomly from }\mathcal{A}_{d},&\quad\text{with probability }\varepsilon\\ \arg\max_{a^{d}\in\mathcal{A}_{d}}Q_{\omega_{1}}\left(s_{t},a^{d},a_{t}^{c}\right),&\quad\text{with probability }1-\varepsilon\end{cases}$
14      Take action $a_{t}=(a_{t}^{d},a_{t}^{c})$, capture the obtained reward $r_{t}$, and transition to the next state $s_{t+1}$;
15
16      Record the transition $(s_{t},a_{t},r_{t},s_{t+1})$ within buffer $\mathcal{B}$;
17
18      Draw a subset of transitions $\{(s_{i},a_{i},r_{i},s_{i+1})\}_{i\in[B]}$ from $\mathcal{B}$;
19
20      Calculate the target continuous action
@@
21      Calculate the target value
@@
22      Perform a single step of gradient descent to update the two Q-networks
$\omega_{j}\leftarrow\omega_{j}-\alpha\nabla_{\omega_{j}}\frac{1}{|B|}\sum_{i\in[B]}\left(Q_{\omega_{j}}\left(s_{i},a_{i}^{d},a_{i}^{c}\right)-y_{i}\right)^{2},\qquad\text{for }j=1,2$
23      if t mod $K$ == 0 then
24             Perform a single step of gradient ascent to update the actor network
$\theta\leftarrow\theta+\beta\nabla_{\theta}\frac{1}{|B|}\sum_{i\in[B]}\max_{a^{d}}Q_{\omega_{1}}\left(s_{i},a^{d},\mu_{\theta}(s_{i})\right)$
25            Adjust the target networks as follows
$\omega_{j}^{\prime}$ $\leftarrow\rho\omega_{j}+(1-\rho)\omega_{j}^{\prime},\qquad\text{for }j=1,2$
$\theta^{\prime}$ $\leftarrow\rho\theta+(1-\rho)\theta^{\prime}$
@@ 25            Adjust the target networks as follows
3 PHEV Energy Management Problem Formulation
@@
3.1 PHEV Powertrain Configuration
@@
Table 1: Technical parameter values of the PHEV.
Symbol Parameters Values
$\omega_{max}$ Maximum shaft angular velocity 250 rad/s
$T_{b,max}$ Maximum mechanical brake torque 6000 Nm
$SOC_{0}$ SOC initial value 0.9
$SOC_{l}$ Minimum SOC 0.3
$SOC_{h}$ Maximum SOC 0.9
$i_{d}$ Final drive gear ratio 4.11
$\eta_{f}$ Final drive efficiency 0.931
$\eta_{g}$ Transmission efficiency 0.931
$Q_{max}$ Battery capacity 26 Ah
$N_{b}$ Number of battery cells 112
$\mu$ Rolling resistance coefficient 0.01
$\rho$ Air density 1.1985 kg/m3
$C_{d}$ Air drag coefficient 0.65
$\theta$ Road grade 0 rad
$A$ Vehicle cross-sectional area 6.73 m2
$m$ Vehicle mass 5000 kg
$r$ Tyre radius 0.5715 m
$\Delta t$ Control interval 1 s
This study focuses on a parallel hybrid electric vehicle [39]. The powertrain architecture comprises an an electric motor, internal combustion engine, an automated manual transmission, and a clutch, illustrated in Fig. 3. The vehicle’s HEV mode can be altered by engaging or disengaging the clutch between the engine and the motor, providing three modes of operation: motor-only, engine-only, and hybrid-drive. The technical parameter values of the vehicle are listed in Table 1.
3.2 Vehicle Dynamics
The equations governing the vehicle dynamics are given as follows:
$T_{w}=\left[ma_{cc}+\mu mg\cos(\theta)+mg\sin(\theta)+\frac{1}{2}AC_{d}\rho v^{2}\right]r$ (15)
@@ 3.2 Vehicle Dynamics
Here, $T_{w}$ represents the torque at the wheels, $m$ denotes the mass of the vehicle, $a_{cc}$ is the acceleration, $\mu$ is the coefficient of rolling resistance, $g$ is the gravitational acceleration, $\theta$ describes the incline of the road, $A$ is the frontal cross-sectional area, $C_{d}$ is the coefficient of drag, $\rho$ stands for air density, $v$ indicates the vehicle’s velocity, $r$ is the radius of the tires, $i_{d}$ refers to the final drive ratio, and $i_{g}$ represents the gear ratio of the transmission, selectable from a set ${\{6.25,3.583,2.22,1.36,1,0.74\}}$.
@@
where $T_{e,d}$ denotes the torque delivered by the engine, and $T_{m}$ is the torque from the motor, both contributing to the shaft torque. $c$ is the clutch state with $c=1$ representing the clutch engaged and $c=0$ representing the clutch disengaged. $\eta_{d}$ and $\eta_{g}$ represent the efficiencies of the final drive and transmission, respectively, and $T_{b}$ is the mechanical brake torque.
3.3 Engine Model
@@
The engine torque consists of two distinct elements: the idle torque $T_{e,i}$ and the driving torque $T_{e,d}$, which can be expressed as:
$T_{e}=T_{e,i}+T_{e,d}$ (18)
@@
where $T_{e,max}=f_{e}(\omega_{e})$ represents the maximum engine torque at a given angular velocity $\omega_{e}$, which is depicted in Fig. 4. When the clutch is disengaged, the torque from the engine, $T_{e,d}$, defaults to zero and $T_{e}$ shifts to $T_{e,i}$, which is primarily utilized to operate the auxiliary systems. The operating speed of the engine is determined by the equation below:
@@
where $\omega_{e}$ represents the engine’s angular velocity, and $\omega_{ei}$ specifies the engine’s idle speed. The constraint on $\omega_{e}$ is imposed to ensure that the engine speed does not drop below the idle speed when the clutch is engaged.
@@ 3.4 Motor Model
3.5 Battery Model
Figure 6: The battery cell’s open-circuit voltage and resistance.
We consider a lithium-ion battery pack consisting of 112 cells, where the cells are connected in series. The battery model is based on the open-circuit voltage $E$ and state of charge (SOC). The SOC is calculated by:
$\dot{SOC}$ $=-\frac{I_{b}}{Q_{max}}$ (28)
$EI_{b}$ $=I_{b}^{2}R_{b}+P_{b}$ (29)
$P_{b}$ $=T_{m}\omega\eta_{m}^{\text{sgn}(-T_{m})}$ (30)
@@ 3.5 Battery Model
where $I_{b}$ is the net current, $Q_{max}$ is the battery capacity, $R_{b}$ is the battery resistance, $P_{b}$ is the battery power, $\eta_{m}$ is the motor efficiency, and $f_{\eta}$ is the motor efficiency interpolation function, which is shown in Fig. 5. The battery open-circuit voltage and resistance are calculated by:
$E$ $=f_{E}(SOC)$ (32)
$R_{b}$ $=f_{R}(SOC).$ (33)
where $f_{E}$ and $f_{R}$ are interpolation functions, as illustrated in Fig. 6. The battery SOC is constrained by:
$SOC_{l}\leq SOC\leq SOC_{h}$ (34)
where $SOC_{l}$ and $SOC_{h}$ are the lower and upper SOC limits, respectively.
3.6 Problem Formulation
Figure 7: CHTC-LT drive cycle for light trucks.
@@ 3.6 Problem Formulation
where $k_{f}$ and $k_{b}$ are the fuel and electricity prices, respectively, $\tilde{p}$ denotes the reference penalty, established at the median value of fuel consumption, $p_{\varsigma}$ and $p_{c}$ are the penalties for gear shift and clutch engagement/disengagement, respectively, $\zeta$ and $\xi$ are the penalty coefficients, $\Delta t$ is the control interval, $T_{0}$ and $T_{cyc}$ are the initial and final time steps of the drive cycle, respectively. The drive cycle utilized for training is the China heavy-duty transient cycle for light trucks (CHTC-LT), depicted in Fig. 7.
4 Experiments and Discussion
4.1 TD3AQ for PHEV Energy Management
4.1.1 State
@@ 4.1.2 Action
where $T_{e}$ is continuous, and $\varsigma$ and $c$ are discrete, making it an MIOC problem. In negative demand torque, regenerative braking is prioritized for battery charging. If regenerative braking is insufficient, mechanical braking provides the remaining torque. When the required torque is positive, the mechanical braking torque is zeroed out, with both the engine and motor contributing to meet the torque requirements. Thus, the proposed action determines the vehicle’s torque distribution entirely.
@@
4.1.3 Reward
In addition to the cost function defined in Eq. (36), it is imperative to ensure that the system dynamics remain within specific constraints. Consequently, the reward function incorporates both fuel usage and penalties for constraint violations. The detailed hand-crafted reward function is defined as follows:
@@ 4.1.3 Reward
4.1.4 Hyperparameters
Table 2: Hyperparameters of TD3AQ.
Parameter Value
Soft target update coefficient 0.001
Reward discount factor 0.99
Optimizer Adam [40]
Actor-network learning rate 0.0001
Q-network learning rate 0.001
Experience replay memory size 200 000
Minibatch size 128
Exploration noise $\mathcal{N}(0,0.02)$
Actor-network hidden layer size $64\times 64$
Q-network hidden layer size $64\times 64$
The hyperparameters of the proposed TD3AQ are summarized in Table 2, both actor and Q networks consist of two hidden layers with 64 neurons each. The continuous action noise is modeled as a zero-mean Gaussian distribution with a standard deviation of 0.02.
4.2 Comparison Benchmarks
Four benchmark methods, DP, DDAQ [28], PATD3 [18], and Rainbow [16], are selected for comparison. Although we implemented Hybrid PPO following the approach described by Neunert et al. [25], it was excluded from the comparison due to its inferior performance, incurring a 29.80% higher cost compared to DP.
DP utilizes the MATLAB toolbox created by Sundström et al. [41]. The continuous engine torque is segmented at intervals of 25 Nm. Empirical experiments have demonstrated that finer discretization does not improve the performance significantly (less than 0.05% improvement in cost with 10 Nm discretization while requiring seven times more computation time). Since Rainbow is designed for discrete action spaces, it adopts the same discretization scheme as DP. All RL algorithms utilize Optuna [42] to tune the hyperparameters.
4.3 Experimental Setup
The training process involves a total of 200 000 timesteps, where each timestep represents one second of driving time. Model evaluation was performed every one episode, with each episode corresponding to one drive cycle. The final model for testing was selected based on the highest return achieved during evaluation. The source code of the algorithm is available at https://github.com/xujinming01/TD3AQ.
4.4 Learning Ability Assessment
Figure 8: The reward curves during training.
The reward curves during training are shown in Fig. 8. It can be observed that, with the exception of PATD3, all algorithms start to converge after 50,000 timesteps. Furthermore, TD3AQ exhibits relatively lower fluctuations after convergence. In contrast, DDAQ and Rainbow encounter notable drops in rewards intermittently. Taken together, these results highlight TD3AQ’s expedited convergence and enhanced stability throughout the training process.
4.5 Optimal Control Performance
@@
(d) Clutch status (1 for engaged, 0 for disengaged)
Figure 9: The hybrid action results from TD3AQ and DP for CHTC-LT.
Table 3: Results when trained and tested on the same drive cycle (CHTC-LT).
@@
TD3AQ 55 136 0 0 0 15.74 4.69%
DDAQ 126 183 0 0 0 16.25 8.06%
PATD3 70 106 1 1 0 16.49 9.66%
Rainbow 171 160 1 0 0 16.12 7.19%
Fig. 9 displays the hybrid action output of DP and TD3AQ employed in CHTC-LT. It is evident that TD3AQ exhibits similar output trends to DP, however, with a higher frequency of gear shifts and clutch state changes. Notably, DP operates under the assumption that the entire drive cycle is known in advance. In contrast, RL algorithms only possess access to the current state, yielding suboptimal solutions. Table 3 presents the simulation results for CHTC-LT, where $\tau_{sh}$ and $\tau_{cl}$ denote the number of gear shifts and clutch engagements, respectively, $\nu_{T_{m}}$, $\nu_{\omega}$, and $\nu_{soc}$ denote the instances where motor torque, shaft angular velocity, and SOC exceed the predefined constraints. Among the RL algorithms, TD3AQ exhibits the lowest cost, with a gap of 4.69% compared to DP. Furthermore, TD3AQ and DDAQ do not violate any constraints, while PATD3 and Rainbow demonstrate a few violations in terms of motor torque and shaft angular velocity constraints.
@@
Table 4: Wall-clock time comparison for CHTC-LT.
Strategy Training time Execution time
TD3AQ 0.21 h 0.46 ms
DDAQ 0.20 h 0.45 ms
PATD3 0.27 h 1.04 ms
Rainbow 0.85 h 1.15 ms
DP N/A 96599 ms
Table 4 presents the time consumption comparison for CHTC-LT, where execution time represents the time required to execute a single action during testing. While DP can obtain the optimal solution for each step in a single calculation, it comes at the expense of a significant amount of time, with its single solution time being six orders of magnitude greater than RL. This substantial discrepancy in time consumption highlights the efficiency of the RL algorithms, as indicated by their step times falling within the millisecond range. Consequently, RL demonstrates promising potential for real-time optimal control applications.
4.6 Generalization Results
Table 5: Generalization results for new drive cycles.
@@ 4.6 Generalization Results
TD3AQ 0 0 0 14.09 3.60%
DDAQ 0 0 0 15.15 11.40%
PATD3 0 0 0 16.68 22.60%
Rainbow 0 0 0 14.59 7.21%
@@
TD3AQ 6 10 0 10.12 8.55%
DDAQ 22 0 0 10.28 10.20%
PATD3 15 125 0 10.81 15.88%
Rainbow 24 5 0 10.05 7.73%
@@
TD3AQ 0 0 0 5.72 3.64%
DDAQ 0 1 0 6.70 21.33%
PATD3 0 8 0 5.74 3.95%
Rainbow 1 0 0 5.94 7.69%
To evaluate the generalization capabilities of the proposed HARL algorithm TD3AQ, the model trained exclusively on the CHTC-LT drive cycle is tested on three additional drive cycles: JE05, HD-UDDS, and WHVC. The simulation results are summarized in Table 5.
The findings demonstrate that the TD3AQ algorithm exhibits commendable generalization performance when applied to new drive cycles. In comparison with other RL algorithms, TD3AQ closely approximates the cost achieved by the DP strategy, with an average gap of 5.26% across the three cycles. However, the remaining RL algorithms display a larger deviation in terms of cost.
@@
5 Conclusion
In this paper, we present a novel reinforcement learning algorithm called TD3AQ for solving mixed-integer optimal control problems. The TD3AQ algorithm uses actor-Q networks to directly generate continuous and discrete actions from the current state. We evaluate the proposed TD3AQ algorithm on a hybrid electric vehicle energy management problem and show that it achieves satisfactory performance with a cost gap of 4.69% compared to the DP strategy. TD3AQ’s ability to generalize is demonstrated through the results on three new drive cycles. The algorithm’s sub-millisecond response time indicates its significant potential for real-time, mixed-integer optimal control applications. However, note that the algorithm carries the risk of violating constraints, which, if unaddressed, could potentially lead to vehicle and battery damage in real-world scenarios. Future work will focus on extending TD3AQ to more complex, higher-dimensional problems to further substantiate its adaptability and robustness, and developing HARL algorithms with safety guarantees.
Acknowledgment
This research received partial funding from the Guangzhou Basic and Applied Basic Research Program under Grant 2023A04J1688, and was also supported by South China University of Technology faculty start-up fund.